OpenStudy (anonymous):

Find the equation of a circle that has its center at (-4, 3) and has a radius of 5.

OpenStudy (anonymous):

these are the answers x + 4) 2 + (y - 3) 2 = 25 (x - 4) 2 + (y - 3) 2 = 25 (x + 4) 2 + (y + 3) 2 = 25 (x - 4) 2 + (y + 3) 2 = 25

OpenStudy (projects):

http://www.mathwarehouse.com/geometry/circle/equation-of-a-circle.php Read through this. It will help you out a lot. c: